# Momentor: Advancing Video Large Language Model with Fine-Grained Temporal Reasoning

### v1

@@ Appendix D Prompts
Hello, and I’d like you to act as a question generator. You will receive two descriptions about one source clip and one target clip. Your task is to generate one question with differences of the two clips to test if a deep learning model can retrieve the target clip based on the content of the source clip by asking about ”could you please find a clip with following differences”. Below are some examples: <Example 1> [Source Clip Content] An old professor is giving a lecture in front of a blackboard. [Target Clip Content] An elderly man is giving a speech in front of a blackboard, holding a ruler. [Major Differences] The man in the target clip content is holding a ruler. [Instruction] Please watch the {{SOURCE_CLIP}}. Could you please find a similar clip, but this time the speaker is holding something at hand? <Example 2> [Source Clip Content] A beautiful scene of primeval forest. [Target Clip Content] A beautiful view of coral reef taken in shallow sea. [Major Differences] The scene in the target clip is a seascape rather than a forest landscape. [Instruction] Please watch the {{SOURCE_CLIP}}. Is there any similar clip with a different kind of scenery? Now given the following descriptions about one source clip and one target clip, please generate one question with differences of the two clips to test if a deep learning model can retrieve the target clip based on the content of the source clip by asking about ”could you please find a clip with following differences”. [Source Clip Content] {source_clip_content} [Target Clip Content] {target_clip_content}

### v2

@@ Appendix D Prompts
Composed Retrieval Hello, and I’d like you to act as a question generator. You will receive two descriptions about one source clip and one target clip. Your task is to generate one question with differences of the two clips to test if a deep learning model can retrieve the target clip based on the content of the source clip by asking about ”could you please find a clip with following differences”. Below are some examples: <Example 1> [Source Clip Content] An old professor is giving a lecture in front of a blackboard. [Target Clip Content] An elderly man is giving a speech in front of a blackboard, holding a ruler. [Major Differences] The man in the target clip content is holding a ruler. [Instruction] Please watch the {{SOURCE_CLIP}}. Could you please find a similar clip, but this time the speaker is holding something at hand? <Example 2> [Source Clip Content] A beautiful scene of primeval forest. [Target Clip Content] A beautiful view of coral reef taken in shallow sea. [Major Differences] The scene in the target clip is a seascape rather than a forest landscape. [Instruction] Please watch the {{SOURCE_CLIP}}. Is there any similar clip with a different kind of scenery? Now given the following descriptions about one source clip and one target clip, please generate one question with differences of the two clips to test if a deep learning model can retrieve the target clip based on the content of the source clip by asking about ”could you please find a clip with following differences”. [Source Clip Content] {source_clip_content} [Target Clip Content] {target_clip_content}